If two events 𝐴 and 𝐵 are independent, and you know that 𝑃(𝐴)=3/10, what is the value of 𝑃(𝐴|𝐵)?

Answer by Edwin McCravy(20054)

If two events are independent then if we are given that one is known to have
occurred, the probability that the other has or will occur is UNCHANGED!

So if we are given that B has occurred (which is what "|B" means), the
probability of A is still 3/10.  So P(A|B) = P(A) = 3/10. 

[Caution: don't get "independent events" confused with "mutually exclusive
events".  Mutually exclusive events cannot both happen.  If one happens, the
other cannot happen.  Independent events CAN both happen, but the
probability of either one is not affected if the other one occurs (or
doesn't occur).]

Edwin